The 𝕓𝕓𝕠𝕝𝕕 symbol font

Alan Jeffrey

April 2002

Introduction

This document describes the 𝕓𝕓𝕠𝕝𝕕 math alphabet. This is an open (or ‘blackboard bold’) geometric sans serif, intended mainly for use in mathematics, but which may prove useful for decorative work.

To use the fonts in L^{A}TEX 2ε, you select the bbold package, and then use the \mathbb command to get blackboard bold mathematics, \textbb for text, or \bbfamily for longer text. The full font is:

´ 0  ´ 1  ´ 2  ´ 3  ´ 4  ´ 5  ´ 6  ´ 7
´ 00x  ℾ  Δ  Θ  Λ  Ξ  ℿ  ⅀  Υ ˝ 0x ´ 01x  Φ  Ψ  Ω  α  β  ℽ  δ  ε
´ 02x  ζ  η  θ  ι  κ  λ  μ  ν ˝ 1x ´ 03x  ξ  ℼ  ρ  σ  τ  υ  ϕ  χ
´ 04x  ψ  !  ”  #  $  %  &  ’ ˝ 2x ´ 05x  ⦅  ⦆  *  +  ,  -  .  /
´ 06x  𝟘  𝟙  𝟚  𝟛  𝟜  𝟝  𝟞  𝟟 ˝ 3x ´ 07x  𝟠  𝟡  :  ;  <  ·  >  ?
´ 10x  @  𝔸  𝔹  ℂ  𝔻  𝔼  𝔽  𝔾 ˝ 4x ´ 11x  ℍ  𝕀  𝕁  𝕂  𝕃  𝕄  ℕ  𝕆
´ 12x  ℙ  ℚ  ℝ  𝕊  𝕋  𝕌  𝕍  𝕎 ˝ 5x ´ 13x  𝕏  𝕐  ℤ  ⟦  \  ⟧  ⟪  ⟫
´ 14x  ‘  𝕒  𝕓  𝕔  𝕕  𝕖  𝕗  𝕘 ˝ 6x ´ 15x  𝕙  𝕚  𝕛  𝕜  𝕝  𝕞  𝕟  𝕠
´ 16x  𝕡  𝕢  𝕣  𝕤  𝕥  𝕦  𝕧  𝕨 ˝ 7x ´ 17x  𝕩  𝕪  𝕫  –  ⫾  —  “  ω

           ˝ 8    ˝ 9    ˝ A    ˝ B    ˝ C    ˝ D    ˝ E    ˝ F

𝕋𝕙𝕖 𝕓𝕓𝕠𝕝𝕕 𝕗𝕠𝕟𝕥 𝕗𝕒𝕞𝕚𝕝𝕪.

This was the first full font I implemented, and has a number of dubious design features. It is very geometric, and the stress does not conform to that of most Latin fonts—this was an attempt to imitate real blackboard handwriting. The difference in widths between ‘𝕒’ and ‘𝕓’ is very noticable. The lip of the ‘𝕣’ is horrible. However, the uppercase latin letters look passable, and they’re almost the only ones anyone ever uses.

Legal rubbish

This package is copyright \begin{matrix}c \\ ◯ \end{matrix} 1989–2002 Alan Jeffrey. All rights are reserved. The moral right of the author has been asserted.

Redistribution and use in source and binary forms, with or without modification, are permitted provided that the following conditions are met:

1. Redistributions of source code must retain the above copyright notice, this list of conditions and the following disclaimer.

2. Redistributions in binary form must reproduce the above copyright notice, this list of conditions and the following disclaimer in the documentation and/or other materials provided with the distribution.

3. The name of the author may not be used to endorse or promote products derived from this software without specific prior written permission.

THIS SOFTWARE IS PROVIDED BY THE AUTHOR “AS IS” AND ANY EXPRESS OR IMPLIED WARRANTIES, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MERCHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE ARE DISCLAIMED. IN NO EVENT SHALL THE AUTHOR BE LIABLE FOR ANY DIRECT, INDIRECT, INCIDENTAL, SPECIAL, EXEMPLARY, OR CONSEQUENTIAL DAMAGES (INCLUDING, BUT NOT LIMITED TO, PROCUREMENT OF SUBSTITUTE GOODS OR SERVICES; LOSS OF USE, DATA, OR PROFITS; OR BUSINESS INTER-RUPTION) HOWEVER CAUSED AND ON ANY THEORY OF LIABILITY, WHETHER IN CONTRACT, STRICT LIABILITY, OR TORT (INCLUDING NEGLIGENCE OR OTHERWISE) ARISING IN ANY WAY OUT OF THE USE OF THIS SOFTWARE, EVEN IF ADVISED OF THE POSSIBILITY OF SUCH DAMAGE.